# Multi-fidelity learning for interatomic potentials: Low-level forces and high-level energies are all you need

### v1

@@ 2.3 Models
A comprehensive set of training tests is performed with each level of theory for both single-fidelity (SF) and multi-fidelity (MF) machine learning interatomic potentials (MLIPs). Table 1 gives a summary of all MLIPs trained in this work, as well as the shorthand notation adopted to distinguish between models. The SF-MLIPs allow for a baseline comparison with the MF-MLIPs as well as a benchmark comparison with the published ANI MLIPs trained to the same level of theory. See Supporting Information for model and training details.
Table 1: All MLIPs trained in this work. SF = single-fidelity. MF = multi-fidelity. DZ = DFT/DZ. TZ = DFT/TZ. MP2 = MP2/TZ. CC = CCSD(T)/CBS. E = energies. F = forces. ✓signifies MLIP was trained with data from this level of theory (DFT/DZ, DFT/TZ, MP2/TZ, or CCSD(T)/CBS) and property (E or F).
DFT/DZ DFT/TZ MP2/TZ CCSD(T)/CBS
@@ 3 Results
The second test case of MFL also utilizes DFT/DZ for the low-level data, but instead utilizes MP2/TZ for the high-level data. In this test case, however, the reduced ANI-1x training dataset does not contain forces for MP2/TZ. Therefore, it is not possible to benchmark the MF-DZ/EF-MP2/E MLIP against an SF-MLIP trained directly to MP2/TZ forces (SF-MP2/EF). The benefit of this second test case is that the COMP6 test set contains MP2/TZ forces for comparison.
The third test case of MFL utilizes DFT/DZ for low-level data and CCSD(T)/CBS for high-level data. CC-level forces are not available for either the ANI-1ccx training dataset or the COMP6 test dataset. Thus, not only are we not able to train a SF-MLIP directly to CC-level forces for comparison (SF-CC/EF), but we also cannot test the quality of the MLIP predictions against CC-level forces. Instead, we compare the SF-MLIP and MF-MLIP performance for predicting CC-level conformer energy differences for the subset of the COMP6 dataset.
Performance for SF and MF models is measured by the mean absolute error (MAE) and root-mean-square error (RMSE) of energies and atomic forces. The RMSEs and MAEs are reported as the average error from an ensemble of eight MLIPs, with the corresponding 95% confidence interval. All errors reported in this work correspond to test errors, i.e., the errors computed on either the heldout dataset or the COMP6 testset. No training or validation errors are reported.
@@ 3.1 Test case one: DFT/DZ and DFT/TZ
Furthermore, MFL with low-level forces achieves similar accuracy as training directly to high-level forces (SF-TZ/EF). Specifically, the MF-DZ/EF-TZ/E force errors (4.16 kcal/mol/Å) are only 18% higher than the SF-TZ/EF force errors (3.52 kcal/mol/Å) and the MF-DZ/EF-TZ/E energy errors (1.49 kcal/mol) are actually 7% lower than the SF-TZ/EF energy errors (1.60 kcal/mol). This performance is quite remarkable considering that MF-DZ/EF-TZ/E was not trained with DFT/TZ forces. Therefore, in the absence of high-level forces, MFL with low-level forces is sufficient for training an MLIP that predicts accurate high-level energies and forces.
@@ 3.2 Test case two: DFT/DZ and MP2/TZ
We now investigate MFL with DFT/DZ and MP2/TZ data. Although the training dataset does not contain any MP2/TZ force data, this test case allows for validation against the MP2/TZ forces in the subset of the COMP6 dataset. In the first test case, the two levels of theory (DFT/DZ and DFT/TZ) are relatively similar and computationally inexpensive. In this test case, however, the MP2/TZ forces are quite different from the DFT/DZ forces (RMSE of 5.9 kcal/mol/Å  when comparing COMP6 DFT/DZ and MP2/TZ forces). The more expensive MP2/TZ forces are also assumed to be more accurate than DFT/TZ, considering MP2 is a higher level of theory and both methods use the same basis set. Therefore, this test case presents a more interesting and practical challenge for MFL than test case one.
@@
Table 3: Comparison of energy and force root-mean-square errors (RMSEs) of MP2/TZ prediction, using single-fidelity (SF) and multi-fidelity (MF) MLIPs trained with MP2/TZ energies with/without DFT/DZ forces and energies. MFL with low-level (DFT/DZ) forces improved the prediction on the high-level (MP2/TZ) energies and forces compared to SF training to high-level energies only. No MP2/TZ forces were used during training of MLIPs. Energy RMSEs are computed relative to the MP2/TZ energies for the in-sample test set. Force RMSEs are computed relative to the MP2/TZ forces in the subset of the COMP6 out-of-sample test set. Error bars represent 95% confidence interval from an ensemble of eight models.
@@
Consistent with the first test case, the results in Table 3 demonstrate that MFL with low-level (DFT/DZ) forces is significantly more accurate than SF training to only high-level (MP2/TZ) energies. Specifically, both the energy and force RMSEs for MF-DZ/EF-MP2/E (1.72 kcal/mol and 3.42 kcal/mol/Å) are approximately a factor of two lower than for SF-MP2/E (2.95 kcal/mol and 8.59 kcal/mol/Å), respectively. Furthermore, the force RMSE for MF-DZ/EF-MP2/E on the subset of the COMP6 dataset (3.42 kcal/mol/Å) is significantly lower than the RMSE between DFT/DZ forces and MP2/TZ forces (5.9 kcal/mol/Å). Thus, MFL is outperforming the accuracy that could be achieved by simply learning the low-level (DFT/DZ) forces.
3.3 Test case three: DFT/DZ and CCSD(T)/CBS
@@
Having verified that MFL quantifiably improves the high-level energies and forces for the first two test cases, we now aim at the most challenging and important test case, namely, MFL with DFT/DZ and gold-standard CCSD(T)/CBS data. However, because CC is prohibitively expensive, no CC-level force data are available for direct validation in either the ANI-1ccx training dataset or the COMP6 test dataset. For this reason, we perform an additional test using CC-level conformer energies in the subset of the COMP6 dataset. The conformer energies are reported as energy differences $(\Delta E)$ between all possible pairs of conformers. Predicting accurate conformer energies is essential for determining the most stable conformer of a molecule. Conformer geometry searches can be quite expensive when performed with QM methods, whereas classical FFs may not be sufficiently accurate to correctly identify the minimum-energy conformer geometry. Thus, an MLIP that can predict accurate CC-level conformer energy differences would be extremely valuable.
@@
Table 4 reports the energy and $\Delta E$ RMSEs for a single-fidelity MLIP trained to CCSD(T)/CBS energies (SF-CC/E), a multi-fidelity MLIP trained to DFT/DZ energies and CCSD(T)/CBS energies (MF-DZ/E-CC/E), and a multi-fidelity MLIP trained to DFT/DZ energies and forces and CCSD(T)/CBS energies (MF-DZ/EF-CC/E). See Supporting Information for additional validation results relevant to the DFT/DZ and CCSD(T)/CBS test case.
@@
Table 4: Comparison of energy and conformer $\Delta E$ root-mean-square errors (RMSEs) of CCSD(T)/CBS prediction, using single-fidelity (SF) and multi-fidelity (MF) MLIPs trained with CCSD(T)/CBS energies with/without DFT/DZ forces and/or energies. MFL with low-level (DFT/DZ) forces improved the prediction on the high-level (CCSD(T)/CBS) energies compared to SF training to high-level energies only. No CCSD(T)/CBS forces were used during training of SF or MF MLIPs. Energy RMSEs are computed relative to the CCSD(T)/CBS energies for the in-sample test set. $\Delta E$ RMSEs are computed relative to the CCSD(T)/CBS conformer energies for the COMP6 out-of-sample test set. Error bars represent 95% confidence interval from an ensemble of eight models.
Energy-RMSE (kcal/mol) $\Delta E$-RMSE (kcal/mol)
SF-CC/E 2.78 $\pm$ 0.05 2.65 $\pm$ 0.04
MF-DZ/E-CC/E 2.76 $\pm$ 0.03 2.60 $\pm$ 0.03
MF-DZ/EF-CC/E 1.52 $\pm$ 0.03 1.60 $\pm$ 0.03
@@
Similar to the results for the previous two test cases, the results in Table 4 demonstrate that MFL with low-level (DFT/DZ) forces is significantly more accurate than SF training to only high-level (CCSD(T)/CBS) energies. Specifically, the MF-DZ/EF-CC/E errors are considerably lower than the SF-CC/E errors on both CCSD(T)/CBS energies (1.52 kcal/mol vs 2.78 kcal/mol) and conformer energy differences (1.60 kcal/mol vs 2.65 kcal/mol), respectively. The MF-DZ/EF-CC/E energy and $\Delta E$ RMSEs (1.52 kcal/mol and 1.60 kcal/mol) are also substantially lower than the RMSEs reported for the ANI-1ccx MLIP (2.57 kcal/mol and 2.07 kcal/mol), respectively. See Supporting Information for additional comparisons between HIP-NN and ANI models. By contrast, MFL without forces (i.e., only energies for both levels of theory) is essentially as accurate as SF training to just the high-level energies, consistent with previous work. [30] This result demonstrates that including low-level forces is essential for MFL to achieve significant improvement in accuracy.
@@ 4 Discussion
These three test cases clearly demonstrate that MFL with low-level forces and high-level energies outperforms SF training with only high-level energies. Moreover, MFL with low-level forces is of similar accuracy as SF training directly to high-level forces, in cases where those are available for comparison. To help quantify the improvement MFL provides in all three test cases, Figure 3 presents the percent decrease in RMSE for energies, forces, and conformer energy differences between each MF-MLIP trained with low-level (DFT/DZ) forces and the corresponding SF-MLIP trained with only high-level (either DFT/TZ, MP2/TZ, or CCSD(T)/CBS) energies. See Supporting Information for additional analysis of percent improvement.
@@
The percent improvement in energy and conformer energy differences with MFL is between 40% and 45% for all three test cases. The percent improvement in forces with MFL is 70% and 60% for MF-DZ/EF-TZ/E tested on the ANI-1x DFT/TZ forces and MF-DZ/EF-MP2/E tested on the COMP6 MP2/TZ forces, respectively. By contrast, the percent improvement with transfer learning (TL) reported for the ANI-1ccx MLIP is considerably smaller (between 15-20% for energies and conformer energies) than with any of the MFL test cases. For a truly fair comparison between TL and MFL, however, future work should investigate the percent improvement for TL that is achieved when pre-training with the same HIP-NN architecture and the same reduced ANI-1x dataset with both DFT/DZ energies and DFT/DZ forces.
One unexpected result is that MFL leads to a measurable decrease in accuracy on the low-level forces themselves (see Supporting Information). Specifically, the RMSEs on the DFT/DZ forces for MF-DZ/EF-TZ/E (3.95 kcal/mol/Å), MF-DZ/EF-MP2/E (3.91 kcal/mol/Å), and MF-DZ/EF-CC/E (3.76 kcal/mol/Å) are all markedly higher than the force RMSE for SF-DZ/EF (3.17 kcal/mol/Å). This increase in error on the low-level (DFT/DZ) forces suggests that the current MFL approach requires a compromise in accuracy on the low-level forces to achieve improved accuracy on the high-level forces. Some possible considerations to help reduce the low-level force errors are to use a larger HIP-NN architecture (i.e., more interactions layers, more atom layers, more features), and/or modify the number of shared parameters, and/or include non-linear layers for the fidelity-specific output nodes. However, we emphasize that the primary goal of this work is to improve the high-level energy and force predictions. Indeed, if only accurate prediction of the low-level energy and forces is the desired outcome, SF learning would suffice.
For completeness, we also investigate the potential benefit of applying MFL with forces from more than one level of theory. In this case, the energies and forces for both DFT/DZ and DFT/TZ are included in training along with the CCSD(T)/CBS energies. Notably, the CC-level errors are actually considerably worse for MFL with three datasets than MFL with just two datasets (see Supporting Information). With twice as much DFT-level data, the worse performance with three datasets is likely attributed to a data imbalance. An alternative explanation is that the DFT/DZ and DFT/TZ forces are providing contradictory information that MFL simply cannot resolve. Another possibility is that the MLIP has reached its capacity to match the different datasets with the chosen architecture and hyperparameters. However, with the current MFL approach, low-level forces from just a single level of theory are sufficient to achieve the desired high-level accuracy.

### v2

@@ 2.3 Models
A comprehensive set of training tests is performed with each level of theory for both single-fidelity (SF) and multi-fidelity (MF) machine learning interatomic potentials (MLIPs). Table 1 gives a summary of all MLIPs trained in this work, as well as the shorthand notation adopted to distinguish between models. The SF-MLIPs allow for a baseline comparison with the MF-MLIPs as well as a benchmark comparison with the published ANI MLIPs trained to the same level of theory. An ensemble consisting of eight MLIPs is trained for each model. See Supporting Information for model and training details.
Table 1: All MLIPs trained in this work. SF = single-fidelity. MF = multi-fidelity. DZ = DFT/DZ. TZ = DFT/TZ. MP2 = MP2/TZ. CC = CCSD(T)*/CBS. E = energies. F = forces. ✓signifies MLIP was trained with data from this level of theory (DFT/DZ, DFT/TZ, MP2/TZ, or CCSD(T)*/CBS) and property (E or F).
DFT/DZ DFT/TZ MP2/TZ CCSD(T)*/CBS
@@ 3 Results
The second test case of MFL also utilizes DFT/DZ for the low-level data, but instead utilizes MP2/TZ for the high-level data. In this test case, however, the reduced ANI-1x training dataset does not contain forces for MP2/TZ. Therefore, it is not possible to benchmark the MF-DZ/EF-MP2/E MLIP against an SF-MLIP trained directly to MP2/TZ forces (SF-MP2/EF). The benefit of this second test case is that the GDB10to13 test set contains MP2/TZ forces for comparison.
The third test case of MFL utilizes DFT/DZ for low-level data and CCSD(T)*/CBS for high-level data. CC-level forces are not available for either the ANI-1ccx training dataset or the GDB10to13 test dataset. Thus, not only are we not able to train a SF-MLIP directly to CC-level forces for comparison (SF-CC/EF), but we also cannot test the quality of the MLIP predictions against CC-level forces. Instead, we compare the SF-MLIP and MF-MLIP performance for predicting CC-level conformer energy differences for the GDB10to13 dataset.
@@ 3.1 Test case one: DFT/DZ and DFT/TZ
Furthermore, MFL with low-level forces achieves similar accuracy as training directly to high-level forces. Specifically, the MF-DZ/EF-TZ/E force errors (4.16 kcal/mol/Å) are only 18% higher than the SF-TZ/EF force errors (3.52 kcal/mol/Å) and the MF-DZ/EF-TZ/E energy errors (1.49 kcal/mol) are actually 7% lower than the SF-TZ/EF energy errors (1.60 kcal/mol). This performance is quite remarkable considering that MF-DZ/EF-TZ/E was not trained with DFT/TZ forces. Therefore, in the absence of high-level forces, MFL with low-level forces is sufficient for training an MLIP that predicts accurate high-level energies and forces.
@@ 3.2 Test case two: DFT/DZ and MP2/TZ
We now investigate MFL with DFT/DZ and MP2/TZ data. Although the training dataset does not contain any MP2/TZ force data, this test case allows for validation against the MP2/TZ forces in the GDB10to13 dataset. In the first test case, the two levels of theory (DFT/DZ and DFT/TZ) are relatively similar and computationally inexpensive. In this test case, however, the MP2/TZ forces are quite different from the DFT/DZ forces (RMSE of 5.9 kcal/mol/Å when comparing DFT/DZ and MP2/TZ forces in GDB10to13). The more expensive MP2/TZ forces are also assumed to be more accurate than DFT/TZ, considering MP2 is a higher level of theory and both methods use a triple zeta basis set. Therefore, this test case presents a more interesting and practical challenge for MFL than test case one.
@@
Table 3: Comparison of energy and force root-mean-square errors (RMSEs) of MP2/TZ prediction, using single-fidelity (SF) and multi-fidelity (MF) MLIPs trained with MP2/TZ energies with/without DFT/DZ forces and energies. MFL with low-level (DFT/DZ) forces improved the prediction on the high-level (MP2/TZ) energies and forces compared to SF training to high-level energies only. No MP2/TZ forces were used during training of MLIPs. Energy RMSEs are computed relative to the MP2/TZ energies for the in-sample test set. Force RMSEs are computed relative to the MP2/TZ forces in the GDB10to13 out-of-sample test set. Error bars represent 95% confidence interval from an ensemble of eight models.
@@
Consistent with the first test case, the results in Table 3 demonstrate that MFL with low-level (DFT/DZ) forces is significantly more accurate than SF training to only high-level (MP2/TZ) energies. Specifically, both the energy and force RMSEs for MF-DZ/EF-MP2/E (1.72 kcal/mol and 3.42 kcal/mol/Å) are approximately a factor of two lower than for SF-MP2/E (2.95 kcal/mol and 8.59 kcal/mol/Å), respectively. Furthermore, the force RMSE for MF-DZ/EF-MP2/E on the GDB10to13 dataset (3.42 kcal/mol/Å) is significantly lower than the RMSE between DFT/DZ forces and MP2/TZ forces (5.9 kcal/mol/Å). Thus, MFL is outperforming the accuracy that could be achieved by simply learning the low-level (DFT/DZ) forces.
3.3 Test case three: DFT/DZ and CCSD(T)*/CBS
@@
Having verified that MFL quantifiably improves the high-level energies and forces for the first two test cases, we now aim at the most challenging and important test case, namely, MFL with DFT/DZ and gold-standard CCSD(T)*/CBS data. Because CC is prohibitively expensive, no CC-level force data are available for direct validation in either the ANI-1ccx training dataset or the GDB10to13 test dataset. However, Supporting Information provides indirect evidence that MFL with low-level forces improves CC-level forces.
We perform two additional tests using CCSD(T)*/CBS energies for conformers from the GDB10to13 dataset and torsion scans from the Sellers et al. dataset. [74] The conformer energies are reported as energy differences $(\Delta E_{\rm conf})$ between all possible pairs of conformers. The torsion energies are reported as energy differences $(\Delta E_{\rm tors})$ relative to the minimum energy for a given torsion scan. Predicting accurate conformer and torsion energies is essential for determining the most stable structure of a molecule. Conformer geometry searches can be quite expensive when performed with QM methods, whereas classical FFs may not be sufficiently accurate to correctly identify the minimum-energy conformer geometry. [77] Thus, an MLIP that can predict accurate CC-level conformer and torsion energy differences would be extremely valuable for real-life applications, such as drug discovery. [78, 79]
@@
Table 4 reports the energy and $\Delta E_{\rm conf}$ RMSEs for a single-fidelity MLIP trained to CCSD(T)*/CBS energies (SF-CC/E), a multi-fidelity MLIP trained to DFT/DZ energies and CCSD(T)*/CBS energies (MF-DZ/E-CC/E), and a multi-fidelity MLIP trained to DFT/DZ energies and forces and CCSD(T)*/CBS energies (MF-DZ/EF-CC/E). See Supporting Information for additional validation results relevant to the DFT/DZ and CCSD(T)*/CBS test case.
@@
Table 4: Comparison of energy, conformer energy differences ($\Delta E_{\rm conf}$), and torsion energy differences ($\Delta E_{\rm tors}$) root-mean-square errors (RMSEs) of CCSD(T)*/CBS prediction, using single-fidelity (SF) and multi-fidelity (MF) MLIPs trained with CCSD(T)*/CBS energies with/without DFT/DZ forces and/or energies. MFL with low-level (DFT/DZ) forces improved the prediction on the high-level (CCSD(T)*/CBS) energies compared to SF training to high-level energies only. No CCSD(T)*/CBS forces were used during training of SF or MF MLIPs. Energy RMSEs are computed relative to the CCSD(T)*/CBS energies for the in-sample test set. $\Delta E_{\rm conf}$ and $\Delta E_{\rm tors}$ RMSEs are computed relative to the CCSD(T)*/CBS energy differences for the GDB10to13 conformer and torsion scan out-of-sample test sets, respectively. Error bars represent 95% confidence interval from an ensemble of eight models.
Energy-RMSE (kcal/mol) $\Delta E_{\rm conf}$-RMSE (kcal/mol) $\Delta E_{\rm tors}$-RMSE
SF-CC/E 2.78 $\pm$ 0.05 2.65 $\pm$ 0.04 0.71 $\pm$ 0.03
MF-DZ/E-CC/E 2.76 $\pm$ 0.03 2.60 $\pm$ 0.03 0.67 $\pm$ 0.04
MF-DZ/EF-CC/E 1.52 $\pm$ 0.03 1.60 $\pm$ 0.03 0.40 $\pm$ 0.02
@@
Similar to the results for the previous two test cases, the results in Table 4 demonstrate that MFL with low-level (DFT/DZ) forces is significantly more accurate than SF training to only high-level (CCSD(T)*/CBS) energies. Specifically, the MF-DZ/EF-CC/E errors are considerably lower than the SF-CC/E errors on CCSD(T)*/CBS energies (1.52 kcal/mol vs 2.78 kcal/mol), CCSD(T)*/CBS conformer energy differences (1.60 kcal/mol vs 2.65 kcal/mol), and CCSD(T)*/CBS torsion scan energy differences (0.40 kcal/mol vs 0.71 kcal/mol), respectively. The MF-DZ/EF-CC/E energy, $\Delta E_{\rm conf}$, and $\Delta E_{\rm tors}$ RMSEs (1.52 kcal/mol, 1.60 kcal/mol, and 0.40 kcal/mol) are also substantially lower than the RMSEs for the ANI-1ccx MLIP (2.57 kcal/mol, 2.62 kcal/mol, and 0.77 kcal/mol), respectively. See Supporting Information for additional comparisons between HIP-NN and ANI models. By contrast, MFL without forces (i.e., only energies for both levels of theory) is essentially as accurate as SF training to just the high-level energies, consistent with previous work. [45] This result demonstrates that including low-level forces is essential for MFL to achieve significant improvement in accuracy.
@@ 4 Discussion
These three test cases clearly demonstrate that MFL with low-level forces and high-level energies outperforms SF training with only high-level energies. Moreover, MFL with low-level forces is of similar accuracy as SF training directly to high-level forces, in cases where those are available for comparison. To help quantify the improvement MFL provides in all three test cases, Figure 3 presents the percent decrease in RMSE for energies, forces, and conformer energy differences between each MF-MLIP trained with low-level (DFT/DZ) forces and the corresponding SF-MLIP trained with only high-level (either DFT/TZ, MP2/TZ, or CCSD(T)*/CBS) energies. See Supporting Information for additional analysis of percent improvement.
@@
The percent improvement in energy and conformer energy differences with MFL is between 40% and 45% for all three test cases. The percent improvement in forces with MFL is 70% and 60% for MF-DZ/EF-TZ/E tested on the ANI-1x DFT/TZ forces and MF-DZ/EF-MP2/E tested on the GDB10to13 MP2/TZ forces, respectively. By contrast, the percent improvement with transfer learning (TL) reported for the ANI-1ccx MLIP is considerably smaller (between 15-20% for energies and conformer energies) than with any of the MFL test cases. For a truly fair comparison between TL and MFL, however, future work could investigate the percent improvement for TL that is achieved when pre-training with the same HIP-NN architecture and the same reduced ANI-1x dataset with both DFT/DZ energies and DFT/DZ forces.
One unexpected result is that MFL leads to a measurable decrease in accuracy on the low-level forces themselves (see Supporting Information). Specifically, the RMSEs on the DFT/DZ forces for MF-DZ/EF-TZ/E (3.95 kcal/mol/Å), MF-DZ/EF-MP2/E (3.91 kcal/mol/Å), and MF-DZ/EF-CC/E (3.76 kcal/mol/Å) are all markedly higher than the force RMSE for SF-DZ/EF (3.17 kcal/mol/Å). This increase in error on the low-level (DFT/DZ) forces suggests that the current MFL approach requires a compromise in accuracy on the low-level forces to achieve improved accuracy on the high-level forces. Future work could attempt to elucidate whether this is a fundamental limitation of MFL, a limitation of our implementation, or a limitation in the model architecture. Some possible considerations to help reduce the low-level force errors are to use a larger HIP-NN architecture (i.e., more interactions layers, more atom layers, more features), and/or modify the number of shared parameters, and/or include non-linear layers for the fidelity-specific output nodes. However, we emphasize that the primary goal of this work is to improve the high-level energy and force predictions. Indeed, if only accurate prediction of the low-level energy and forces is the desired outcome, SF learning would suffice.
For completeness, we also investigate the potential benefit of applying MFL with forces from more than one level of theory. In this case, the energies and forces for both DFT/DZ and DFT/TZ are included in training along with the CCSD(T)*/CBS energies. Notably, the CC-level errors are actually considerably worse for MFL with three datasets than MFL with just two datasets (see Supporting Information). With twice as much DFT-level data, the worse performance with three datasets is likely attributed to a data imbalance. An alternative explanation is that the DFT/DZ and DFT/TZ forces are providing contradictory information that MFL simply cannot resolve. Another possibility is that the MLIP has reached its capacity to match the different datasets with the chosen architecture and hyperparameters. However, with the current MFL approach, low-level forces from just a single level of theory are sufficient to achieve the desired accuracy on the high-level forces.